Solve the following two trigonometric equations:
1. cos 2pix = 1
2. cos 2pix = -1
(Please show your step-by-step solution)

Answer by Alan3354(69443)

1. cos 2pix = 1
cos(2pi*x) = 1
2pi*x = 0 + 2n*pi --- n = integer
x = 0 + n --- n = integer
================
2. cos(2pi*x) = -1
2pi*x = pi + 2n*pi --- n = integer
x = 1/2 + n --- n = integer